OpenStudy (anonymous):

limit as x approaches 0 of 1-cos x/x^2

OpenStudy (anonymous):

\[1-\cos x=2\sin ^{2}\frac{ x }{ 2 }\] make the denominator \[\left( \frac{ x }{2 } \right)^{2}or \frac{ x ^{2} }{ 4 } by multiplying and dividing the denominator by 4 \] Then calculate the limit.

hartnn (hartnn):

or another way is to multiply and divide by (1+cos x)

hartnn (hartnn):

then use the fact that sin^2 x+cos^2 x=1